# Large-Scale Training of Graph Neural Networks for Optimal Markov-Chain Partitioning Using the Kemeny Constant

### v2

@@ 2.3 Graph Neural Networks
To describe GNNs in this work we adopt a message-passing formalism [47] where a GNN-based encoder maps a node $v_{i}$ in ${G}$ to a context-aware embedding using $T$ separate layers via the iterative process:
$\begin{split}m^{t+1}_{i}&=\sum_{j\in\mathcal{N}(v_{i})}M_{t}(h^{t}_{i},h^{t}_{j},x_{ij})\\ h^{t+1}_{i}&=U_{t}(h^{t}_{i},m_{i}^{t+1})\end{split}$ (15)
Where $t\in\{1,...,T\}$ denotes the current network depth, $\mathcal{N}(v_{i})$ is the local neighbourhood of all nodes connected by edges to node $v_{i}$, and $h^{t}_{i}$ represents the hidden representation of node $v_{i}$ at depth $t$. For consistency, the initial node feature vectors $F_{i}$ are represented by $h^{0}_{i}$.
@@
Two different graph layers are tested in our methodology: the GraphSAGE[48] and the GATv2[49] layers. The former was chosen to provide a simpler, lower parameter function, which takes fewer resources to train, while the latter provides a more complex layer with the capability to better attune to all nodes. GraphSAGE layers use a simple message-passing function consisting of either the sum or mean, and a vertex update function which concatenates these expressions and feeds them into a feed-forward neural network.
$m_{i}^{t+1}=\begin{cases}\frac{1}{n}\sum_{j\in\mathcal{N}(v_{i})}h^{t-1}_{j},&\text{if mean}\\ \underset{j\in\mathcal{N}(v_{i})}{\text{max}}\sigma\bigg{(}\mathbf{W}_{\text{pool}}h^{t-1}_{j}+b\bigg{)}&\text{if max}\end{cases}$ (16)
$h_{i}^{t+1}=\sigma(\mathbf{W^{k}}[h_{i}^{t+1}||m_{i}^{t+1}])$ (17)
@@
$m({h}^{t}_{i},{h}^{t}_{j})$ $=\underset{j\in\mathcal{N}(v_{i})}{\text{softmax}}\big{(}\mathbf{a}^{T}\text{LeakyReLU}(\mathbf{W}\cdot[h^{t}_{i}||h^{t}_{j}])\big{)}$ (18)
$h^{t+1}_{i}$ $=\sigma(\sum_{j\in\mathcal{N}(v_{i})}m(h^{t}_{i},h^{t}_{j})\cdot\mathbf{W}h^{t}_{j}),$ (19)
@@ 3 Methodology
The proposed method assumes a given neural network architecture, $\Phi(\theta)$, is sufficiently expressive such that the mappings it explores, from the space of $\mathbf{F}$ to the space of possible partitionings $\mathbf{S}^{N\times M}$, includes the optimal solution to the GP problem given our clustering criteria $\textbf{S}^{\text{opt}}$, along with a region of parameter space which can be optimised via gradient descent to yield it. Taking inspiration from Monte Carlo approaches to similar problems, we look to determine $\textbf{S}^{\text{opt}}$ by initialising $\Phi$ with several different sets of parameters $\theta$ and optimizing them in parallel using modern ML tools. In this framework, GNNs are an effective Ansatz to begin partitioning graphs using complex criteria. They also provide a generic and intuitive optimization framework with gradient descent. As $\Phi$ is a general model, operating in $\mathbb{R}$ without any problem-specific restrictions on its values or gradients, it can yield the trivial solutions to Eq. (14), such as the aforementioned $M=N$ case. To overcome this, we take inspiration from similar work on producing balanced clusters [33, 34], and additionally optimise including a penalty function, which pushes the network away from trivial solutions and gives the final form of the loss function, $\mathcal{L}$:
@@
$\mathcal{L}=\underbrace{\sum_{M}\frac{1}{P^{\text{eq}}_{M}}\sum_{i,j\in M}p^{\text{eq}}_{j}p^{\text{eq}}_{i}t_{ji}}_{\Delta K}+\underbrace{\sum_{j}\frac{\eta}{\sqrt{2\pi\sigma}}\exp{\left[-0.5\left(\frac{\sum_{i}S_{ij}}{\sigma}\right)^{2}\right]}}_{\text{Penalty Function}}$ (20)
@@
Optimal Kemeny constant partitionings can have unbalanced subsets, with clusters that contain only a few nodes often being the most interesting transition states[11]. Therefore the form of the penalty function is designed to penalise for any invalid clustering (where any cluster has zero nodes), while also allowing the network to converge very unbalanced partitions with at least one node. The final penalty function is based on a Gaussian centred on 0, parameterised using $\sigma=1.0$ and $\eta=15$ in this work, meaning the penalty becomes larger as any cluster tends to 0 nodes and is negligible for 1 or more nodes.
@@
As MCs do not fit into the graph structure outlined in Eq. (2), simple methods were derived for interpreting the matrices $\mathbf{A}$ and $\mathbf{F}$ from MCs. The connectivity of the graph was decided using a naive cutoff method, taking all node pairs $(v_{i},v_{j})$ with $(Q_{ij}+Q_{ji})/2>c$, where $c$ is a constant. In practice, we found $c=0.015$ to be sufficient for all graphs we worked with, although in other applications this could be changed depending on the density of the resulting network. A systematic exploration of various node features was undertaken. We tested the adjacency matrix, the rate matrix, the Markov matrix, and the MFPTs, along with a principle component analysis (PCA) of all 3 of these, and the eigenvectors of the rate or Markov matrix which defines the network. We also tested both a simple linear layer which combined them, and finally, a trainable embedding layer, which is often used in dictionaries in large language models [50].
@@
All neural network architectures tested used a GNN-based encoder, mapping the $d_{\text{feat}}$ dimensional features into an embedding dimension $d_{\text{embed}}$ via a hidden representation of size $d_{\text{hidden}}$. Mean pooling was used for our aggregation function in the GraphSAGE layers as it provided better stability during training. The encoder was then followed by either a series of linear layers or attention layers of dimension $d_{\text{hidden}}$ as a decoder [51]. The transformer-like attention-based decoder can process a whole clustering at once as a sequence of node embeddings and was chosen to provide another comparison between a simpler model and one which is larger, more expressive, and better suited to deal with highly contextual sequences like the node-cluster assignment matrix. The output of the transformer layers was normalised and added back to the input before being processed by a feed-forward layer, and we used two of these layers with a dropout layer between. For positional encoding, a single trainable graph layer was used to provide global context for the nodes, and allow the network to better discriminate between nodes with similar features. All neural networks used are described in detail in Fig. (1).
@@
Figure 1: Details of all four GNN architectures used, along with a description of the preprocessing required to begin training, together with an overview of the training loop process for a single gradient optimization step. Implementations of the GraphSAGE and GATv2 GNN layers used in the encoder is provided, along with a schematic description of the transformer decoder (TD) and linear decoder (LD) architectures. Activation functions for the various layers are not included for brevity.
@@
To enable fast and efficient training of the networks, a general parallel methodology for training GNNs for partitioning was established, outlined in Fig. (2). This requires three additional functions to be defined, as well as a valid loss function and the number of processes $P$:
Generator Function $\mathcal{I}$: This function returns valid parameters $\theta$ for a given neural network $\Phi$. This provides the initial starting point in parameter space for optimisation, and complex generators can be used to provide a quicker convergence. In this work, however, simple initialisation using both the Kaiming and Glorot algorithms were tested, before settling on using Glorot due to its slight advantage observed in early tests.[52, 53]
Culling Function $\mathcal{C}$: This function determines whether or not a worker thread has finished training a network, and is called after a predetermined number of optimisation steps has finished. Observing individual runs, it is common to see neural networks which initialise in local minima of parameter space and are unable to be trained. An effective culling function allows for these types of runs to be terminated early, allowing for more time to be spent in interesting regions of parameter space. To show the validity of the method, all experiments in this paper terminate after a predetermined number of training steps.
@@
To test the validity and effectiveness of using GNNs to cluster Markov chains, the method was first tested on synthetic kinetic graphs. Firstly, a simple stochastic block model (SBM) [54] was employed to generate random networks with 4 neighbourhoods. The initial set of networks used sampled inter-cluster connection probabilities from a Gaussian distribution scaled to be in the range $[0,0.25]$ and set intra-cluster probabilities to a constant 0.5. This led to networks with dense clusters, where the optimal partitioning always successfully separated them. We also designed SBM networks where the optimal clustering was less clear, sampling all connection probabilities from the same Gaussian as before. For both sets of examples, SBM networks with disconnected nodes were discarded, and the random-walk normalized Laplacian matrix, $\mathbf{L}$, was used to generate rate matrices via:
@@
Where $\mathbf{I}$ is the identity matrix, $\mathbf{D}$ is a matrix with the degrees of each node along the leading diagonal, and $\mathbf{A}$ is the adjacency matrix. Results on these synthetic graphs were compared to clusterings and respective $\Delta K$ values obtained from the PCCA+ method [55]. Each SBM experiment trained 100 independently initialised models for 250 steps of gradient descent on 500 randomly generated graphs, ranging from $N=150$ to $N=300$ nodes in increments of 50.
@@
To establish a baseline performance for all models on a single graph, the network was tested on a graph corresponding to a rate matrix derived from an analytical 1D potential simulation[29]. The system consisted of 100 nodes distributed along the potential, which was comprised of 4 clear wells: 3 with 20 nodes and one with 40, with the optimal KC partitioning successfully separating all 4 into individual partitions. This example of a linear chain network architecture was chosen to test the network in situations where connectivity is limited. All 4 models were tested 1000 times on this example for 500 gradient descent steps each. A feature comparison was also performed on this system, where 1000 GraphSAGE + LD models were trained for 500 steps using each set of features described above.
@@
Ultimately, the method was also tested on a network derived from MD data. We used the MSM benchmark system of PyEMMA[56] corresponding to 500 ns long implicit water MD simulations of a small 5-amino-acid system, which we refer to as the pentapeptide system. The simulation was first clustered into 250 states using K-means on tICA projected data, and then a Markov matrix was calculated [57, 58], which was used as the input to our method. Each experiment consisted of training 1000 models for 500 gradient descent steps to cluster the 250-state MSM into 5 states. As this example is the closest to examples derived from biomolecular simulations, we performed an additional hyper-parameter search for the optimal network architecture. For this, we again trained 1000 GraphSAGE + LD models for 500 steps while varying the number of layers in both the decoder and the encoder, along with $d_{\text{hidden}}$ and $d_{\text{embed}}$. Additionally, we again performed a feature comparison in the same way as described earlier. We compared the results of this example to the PCCA+ and PVTC, our previous MCMC approach.[29]
@@
The neural networks and relevant code were implemented in Python 3.10.6 using PyTorch 2.0[59] or Tensorflow. [60] Network diagrams were made using the Netgraph library. [61] To perform gradient descent steps the Adam optimizer[62] was used with a learning rate of $1e^{-4}$. All networks used $d_{\text{embed}}=32$ and $d_{\text{hidden}}=64$, with 4 GNN layers in the encoder and 2 layers in the decoder, unless otherwise specified. Due to the numerical sensitivity and dependence of the MFPT calculation on accurate eigenvalues, 64-bit floats were used for all calculations and they were run on the CPU. Each experiment involved the simultaneous execution of 90 processes in parallel, with each process running on a single core.
@@ 4 Results & Discussion
Our GNN approach successfully found the optimal clustering for all SBM networks in the first set, where the clusters were well-defined, in 100 training runs or less. This matched the performance of the PCCA+ algorithm, confirming the viability of our method in trivial situations. Fig. 3 details the performance on the second set of SBM graphs featuring a more diverse range of networks. Our implementation identified partitionings which better optimised $\Delta K$ than the PCCA+. However, the improvement diminished as the number of nodes in the network increased, as often the clustering became clearer. We found that PCCA+ sometimes returned graphs without the specified number of clusters, and sometimes failed to find an effective partitioning when given unclear clusters. The stochastic nature of the GNN-based method however lends itself more to these systems where several viable partitionings have a low $\Delta K$ value. In these examples the optimal clustering is close to many other similar solutions, therefore it is helpful to evaluate many initialisations to evaluate most of the very low $\Delta K$ partitionings.
@@
Figure 3: The performance of the SBM network clustering for randomly sampled cluster connection probabilities. (a) Example of a partitioning produced by the GNN optimisation. (b) and (c) Quantitative assessment of the method’s performance over 100 runs for the 500 different networks. (b) The proportion of runs where the method yields a partitioning with a lower $\Delta K$ value than PCCA+. (c) The average time needed in seconds for one training run when limited to run on one 2.20GHz CPU core.
@@
The 1D potential (Fig. 4a) was used to test our approach on sparse graphs where connectivity becomes limited, and the majority of effective $\Delta K$ partitionings lie in a small region of partitioning space. In this regime gradient descent steps can become less effective as the majority of initial partitions are far from the region of interest, and can be unable to escape their local minima. Fortunately, simply training more neural networks can alleviate the problem by providing a wider range of starting points for optimization. In future applications, an effective choice of the initialization function $\mathcal{I}$ could speed up convergence, whereby $\mathcal{I}$ can learn and then avoid initializing in uninteresting regions of parameter space. To further test whether different node features could improve the successful optimisation outcomes, we tested several additional node features, sampling the optimisation outcome 1000 times each. In total, we tested nine different node feature types: (i) adjacency matrix, (ii) the PCA vectors created from the adjacency matrix, (iii) rate matrix, (iv) the PCA vectors created from the rate matrix, (v) MFPTs (default in other examples), (vi) the PCA vectors created from the MFPT matrix, (vii) eigenvectors of the rate matrix, (viii) trainable embeddings, (ix) a neural network. The feature comparison results (Fig. 4b) show that it is possible to use improved node features, such as the adjacency matrix or a trainable embedding in this case, which achieve a much more successful optimisation for the 1D model potential-derived network.
@@
Optimal partitionings are highly dependent on the connectivity of the graph. It becomes difficult to discriminate nodes at high energies (with low populations) based solely on the rate matrix, as these nodes have similar transitions. The adjacency matrix, however, provides more information on the node location in the global structure of the network, leading to a noticeable advantage. Our GNN implementation allows for simple changes to define the node features on a problem-specific basis, therefore giving the method additional viability across a wide range of MCs.
@@
Figure 4: (a) 1D potential used for an example of a linear chain network. Nodes are coloured according to the most optimal 4-state clustering for the system. (b) Violin plots for feature comparisons on the 1D potential derived network. Each experiment was run 1000 times with each set of features. The corresponding distribution of final $\Delta K$ values is shown, along with a dotted blue line signifying the $\Delta K$ value of the optimal solution for each problem.
@@
We also tested the four GNN architectures (Fig. 5) training 1000 models of each variety on the 1D potential, using the adjacency matrix as node features. We found that simpler models perform better at finding optimal partitionings over more complex models in this framework. Despite the better average performance larger models provided, none of them were able to find the optimal clustering over 1000 runs. The optimal clustering, however, is found consistently by both by the PVTC and the PCCA+ methods. The wide variance of the observed results showcases the poor ability of all models to optimize initial partitionings on this network when started from many regions of parameter space. The large magnitude of some solutions is indicative of a partitioning problem where there is only a small region of interest in partitioning space and larger models like the GATv2 and TD models struggle to find this parameter space, possibly due to the more complex and unclear gradients calculated when the node embeddings are more co-dependent and additional resources and required to train the extra parameters.
@@
Figure 5: The distribution of $\Delta K$ values from training 1000 copies of each neural network architecture on the 1D potential derived network. A logarithmic scale is used to highlight the large differences between the different models. The blue dashed line indicates the value associated with the optimal clustering for this system.
@@
Table 1: Network archtecture comparison for the 1D potential derived network. The four network architectures are compared in terms of the mean (standard deviation) and minimum for 1000 experiments, along with the percentage of times the optimal solution was found.

### v3

@@ 2.3 Graph Neural Networks
To describe GNNs in this work we adopt a message-passing formalism 49 where a GNN-based encoder maps a node $v_{i}$ in ${G}$ to a context-aware embedding using $T$ separate layers via the iterative process:
$\begin{split}m^{t+1}_{i}&=\sum_{j\in\mathcal{N}(v_{i})}M_{t}(h^{t}_{i},h^{t}_{j},e_{ij})\\ h^{t+1}_{i}&=U_{t}(h^{t}_{i},m_{i}^{t+1})\end{split}$ (18)
where $t\in\{1,...,T\}$ denotes the current network depth, $\mathcal{N}(v_{i})$ is the local neighbourhood of all nodes connected to node $v_{i}$ by edges $e_{ij}$, and $h^{t}_{i}$ represents the hidden representation of node $v_{i}$ at depth $t$. For consistency, the initial node feature vectors $F_{i}$ are represented by $h^{0}_{i}$.
@@
Two different graph layers are tested in our methodology: the GraphSAGE50 and the GATv251 layers. The former was chosen to provide a simpler, lower parameter function, which takes fewer resources to train, while the latter provides a more complex layer with the capability to better attune to all nodes. GraphSAGE layers use a simple message-passing function consisting of either the sum or mean, and a vertex update function which concatenates these expressions and feeds them into a feed-forward neural network:
$m_{i}^{t+1}=\begin{cases}\frac{1}{n}\sum_{j\in\mathcal{N}(v_{i})}h^{t}_{j},&\text{if mean}\\ \underset{j\in\mathcal{N}(v_{i})}{\text{max}}\sigma\bigg{(}\mathbf{W}_{\text{pool}}h^{t}_{j}+b\bigg{)}&\text{if max}\end{cases}$ (19)
$h_{i}^{t+1}=\sigma\Bigg{(}\mathbf{W^{k}}[h_{i}^{t}||m_{i}^{t+1}]\Bigg{)}$ (20)
@@
$m({h}^{t}_{i},{h}^{t}_{j})$ $=\underset{j\in\mathcal{N}(v_{i})}{\text{softmax}}\Bigg{(}\mathbf{a}^{T}\text{LeakyReLU}(\mathbf{W}\cdot[h^{t}_{i}||h^{t}_{j}])\Bigg{)}$ (21)
$h^{t+1}_{i}$ $=\sigma\Big{(}\sum_{j\in\mathcal{N}(v_{i})}m(h^{t}_{i},h^{t}_{j})\cdot\mathbf{W}h^{t}_{j}\Big{)}$ (22)
@@ 3 Methodology
The proposed method assumes that a given neural network architecture, $\Phi(\theta)$, is sufficiently expressive such that the mappings it explores, from the space of $d_{\text{feat}}$ dimensional node features, $\mathbf{F}\in\mathcal{R}^{N\times d_{\text{feat}}}$, to the space of possible $M$-state partitionings, $\mathbf{S}\in\mathcal{R}^{N\times M}$, includes the optimal solution, $\textbf{S}^{\text{opt}}$, to the GP problem given our clustering criteria, along with a region of parameter space which can be optimized via gradient descent to yield it. Taking inspiration from Monte Carlo approaches to similar problems, we look to determine $\textbf{S}^{\text{opt}}$ by initializing $\Phi$ with several different sets of parameters $\theta$ and optimizing them in parallel using modern ML tools. In this framework, GNNs are an effective Ansatz to begin partitioning graphs using complex criteria. They also provide a generic and intuitive optimization framework with gradient descent. As $\Phi$ is a general model, operating in $\mathbb{R}$ without any problem-specific restrictions on its values or gradients, it can yield the trivial solutions to Eq. (17), such as the aforementioned $M=N$ case. To overcome this, we take inspiration from similar work on producing balanced clusters 33, 34, and added a penalty function to $\Delta K$, which drives the network away from trivial solutions and gives the final form of the loss function, $\mathcal{L}$:
@@
$\mathcal{L}=\underbrace{\sum_{L}^{M}\frac{1}{P^{\text{eq}}_{L}}\sum_{i,j}^{N}p^{\text{eq}}_{j}p^{\text{eq}}_{i}t_{ji}S_{iL}S_{jL}}_{\Delta K}+\underbrace{\sum_{L}^{M}\frac{\eta}{\sqrt{2\pi\sigma}}\exp{\left[-0.5\left(\frac{\sum_{i}^{N}S_{iL}}{\sigma}\right)^{2}\right]}}_{\text{Penalty Function}}$ (23)
@@
Optimal KC partitionings can have unbalanced subsets, with clusters that contain only a few nodes often being the most interesting transition states11. Therefore, the form of the penalty function is designed to penalise any invalid clustering (where any cluster has zero nodes), while also allowing the network to converge significantly unbalanced partitions with at least one node. The final penalty function is based on a Gaussian centered at 0, parameterized in this work using $\sigma=1.0$ and $\eta=15$. This means the penalty increases as any cluster approaches 0 nodes and becomes negligible for clusters with 1 or more nodes.
@@
As MCs do not fit into the graph structure outlined in Eq. (2), simple methods were derived for interpreting node connectivity using an adjacency matrix, $\mathbf{A}$, and node feature matrix, $\mathbf{F}$. The connectivity of the graph was decided using a cutoff method, taking all node pairs $(v_{i},v_{j})$ with $(Q_{ij}+Q_{ji})/2>c$, where $c$ is a constant. In practice, we found $c=0.015$ to be sufficient for all graphs we worked with, although in other applications this could be changed depending on the density of the resulting network.
A systematic exploration of various node features was undertaken. We tested using the columns of various matrices which can be identified with nodes, such as the adjacency matrix, the rate matrix, the Markov matrix, and the MFPTs, as node features. For each node, we represented its connectivity, transition rates/probabilities to other nodes, or average time required to reach each node, by splitting these matrices into vectors. Alongside these, principal component analysis (PCA), implemented as the singular value decomposition (SVD) of each of these matrices, of all previously mentioned matrices, and the eigenvectors of the rate or Markov matrix were also tested. Finally, we also tested both a simple linear layer, which combined the adjacency, rate/Markov, and MFPT columns for each node, and a trainable embedding layer, similar to those often used as dictionaries in large language models 52. More details regarding the features used are given in SI(A).
@@
All neural network architectures tested used a GNN-based encoder, mapping the $d_{\text{feat}}$ dimensional features into an embedding dimension $d_{\text{embed}}$ via a hidden representation of size $d_{\text{hidden}}$. Mean pooling was used for the aggregation function in the GraphSAGE layers as it provided better stability during training. The encoder was then followed by either a series of linear layers or dot-product attention-based transformer layers of dimension $d_{\text{hidden}}$ as a decoder 53. When using attention, the output of each transformer layer was normalised and added back to the input before being processed by a feed-forward layer. Our implementation used two transformer layers with a dropout layer between them. The output of a single trainable graph layer was used as positional encoding before the decoder. This transformer-like, attention-based decoder can process a whole clustering at once as a sequence of node embeddings. It was chosen to provide a comparison between a simpler model and one that is larger, more expressive, and better suited to handle highly contextual sequences, such as the node-cluster assignment matrix. All neural networks used are described in detail in Fig. 1.
@@
Figure 1: Details of all four GNN architectures used, along with a description of the preprocessing required to begin training, together with an overview of the training loop process for a single gradient optimization step. Implementations of the GraphSAGE and GATv2 GNN layers used in the encoder are provided, along with a schematic description of the transformer decoder (TD) and linear decoder (LD) architectures. Activation functions for the various layers are not included for brevity.
@@
To enable fast and efficient training of the networks, a general parallel methodology for training GNNs for partitioning was established, outlined in Fig. 2. This requires three additional functions to be defined, as well as a valid loss function and the number of processes $P$:
Generator Function $\mathcal{I}$: This function returns valid parameters $\theta$ for a given neural network $\Phi$. This provides the initial starting point in parameter space for optimization, and complex generators can be used to provide a quicker convergence. In this work, however, simple initialization using both Kaiming and Glorot algorithms was tested. We used the Glorot method due to its slight advantage observed in early tests.54, 55 Additionally, we also explored bespoke initializations based on converged models, where a loss function was used that corresponded to the PCCA+ clustering.
Culling Function $\mathcal{C}$: This function determines whether or not a worker thread has finished training a network, and is called after a predetermined number of optimization steps has finished. Observing individual runs, it is common to see neural networks which initialize in local minima of parameter space and are unable to be trained. An effective culling function allows for these types of runs to be terminated early, allowing for more time to be spent in interesting regions of parameter space. To show the validity of the method, all experiments in this paper terminate after a predetermined number of training steps.
@@
To test the validity and effectiveness of using GNNs to cluster Markov chains, the method was first tested on synthetic kinetic graphs. Firstly, a simple stochastic block model (SBM) 56 was employed to generate random networks with 4 neighbourhoods. The initial set of networks used sampled inter-cluster connection probabilities from a Gaussian distribution scaled to be in the range $[0,0.25]$ and set intra-cluster probabilities to a constant 0.5. This led to networks with dense clusters, where the optimal partitioning always successfully separated them. We also designed SBM networks where the optimal clustering was less clear, sampling all connection probabilities from the same Gaussian as before. For both sets of examples, SBM networks with disconnected nodes were discarded, and the random-walk normalized Laplacian matrix, $\mathbf{L}$, was used to generate rate matrices via:
@@
where $\mathbf{I}$ is the identity matrix, $\mathbf{D}$ is a matrix with the degrees of each node along the leading diagonal, and $\mathbf{A}$ is the adjacency matrix given by Eq. (1). Results on these synthetic graphs were compared with clusterings and respective $\Delta K$ values obtained from the PCCA+ clustering. We note that as PCCA+ uses a different kinetics-based objective function29 that reduced the clustering problem to $M^{2}$ dimensions, the optimization is readily solvable using Schur decomposition, therefore only a single optimal cluster is obtained using PCCA+. Each SBM experiment trained 100 independently initialized models for 250 steps of gradient descent on 500 randomly generated graphs, ranging from $N=150$ to $N=300$ nodes in increments of 50.
@@
To establish a baseline performance for all models on a single graph, the network was tested on a graph corresponding to a rate matrix derived from an analytical 1D potential simulation29. The system consisted of 100 nodes distributed along the potential, which comprised four distinct wells: three with 20 nodes, and one with 40 nodes. The optimal KC partitioning successfully separated all 4 wells into individual partitions. This example of a linear chain network architecture was chosen to test the network in situations where connectivity is limited. All four models were tested 1000 times on this example for 500 gradient descent steps each. A feature comparison was also performed on this system, where 1000 GraphSAGE + LD models were trained for 500 steps using each set of features described above.
@@
Ultimately, the method was tested on a network derived from MD data. We used the MSM benchmark system of PyEMMA57 corresponding to 500 ns long implicit water MD simulations of a small 5-amino-acid system, here referred to as the pentapeptide system. The simulation initially underwent clustering into 250 states using K-means on tICA projected data, followed by the calculation of a Markov matrix 58, 59, which served as the input to our method. Each experiment consisted of training 1000 models for 500 gradient descent steps to cluster the 250-state MSM into 5 states. Given the relevance of this example to biomolecular simulations, an additional hyper-parameter search was conducted to identify the optimal network architecture. For this purpose, we again trained 1000 GraphSAGE + LD models for 500 steps while varying the number of layers in both the decoder and the encoder, along with the dimensions $d_{\text{hidden}}$ and $d_{\text{embed}}$. Furthermore, we performed a feature comparison following the same methodology as previously described. We also compared the results of this example to the PCCA+ and parallel tempering variational clustering (PTVC), our previous MCMC approach.29 Additionally, we used the DBI score with the Markov matrix as the node features vectors as a metric to assess the quality of the partitionings produced.
@@
The neural networks and relevant code were implemented in Python 3.10.6 using PyTorch 2.060 or Tensorflow. 61 Network diagrams were generated using the Netgraph library. 62 To perform gradient descent steps, the Adam optimizer63 was employed with a learning rate of $1\times 10^{-4}$. All networks utilized $d_{\text{embed}}=32$ and $d_{\text{hidden}}=64$, with 4 GNN layers in the encoder and 2 layers in the decoder, unless otherwise specified. Due to the numerical sensitivity and dependence of the MFPT calculation on accurate eigenvalues, 64-bit floats were used for all calculations, which were run on the CPU. Each experiment involved the simultaneous execution of 90 processes in parallel, with each process running on a single core. To compute the DBI scores, the scikit-learn implementation was utilized64.
@@ 4 Results & Discussion
Our GNN approach successfully found the optimal clustering for all SBM networks in the first set (e.g., Fig. 3a), where the clusters were well-defined, in 100 training runs or less. The optimal cluster matched the PCCA+ solution, confirming the viability of KC as the optimization criterium. Figs. 3b and 3c detail the performance on the second set of SBM graphs featuring a more diverse range of networks. Our implementation identified partitionings which better optimized $\Delta K$ than the PCCA+ (Fig. S1). However, the improvement diminished as the number of nodes in the network decreased, as often the clustering became clearer. The stochastic nature of the GNN-based method however lends itself more to these systems where several viable partitionings have a low $\Delta K$ value. In these examples the optimal clustering is close to many other similar solutions, therefore it is helpful to evaluate many initializations to evaluate most of the very low $\Delta K$ partitionings. Violin plot distributions of optimal KC values found are also shown in Fig. S1.
@@
Figure 3: The performance of the SBM network clustering for randomly sampled cluster connection probabilities. (a) Example of a partitioning produced by the GNN optimization. (b) and (c) Quantitative assessment of the method’s performance over 100 runs for the 500 different networks. (b) The proportion of runs where the method yields a partitioning with a lower $\Delta K$ value than PCCA+. (c) The average time needed in seconds for one training run when limited to run on one 2.20GHz CPU core.
@@
The 1D potential (Fig. 4a) was used to test our approach on sparse graphs where connectivity is limited to a linear chain, and the majority of effective $\Delta K$ partitionings lie in a small region of partitioning space. In this regime gradient descent steps can become less effective as the majority of initial partitions are far from the region of interest, and can be unable to escape their local minima. Fortunately, simply training more neural networks can alleviate the problem by providing a wider range of starting points for optimization. In future applications, an effective choice of the initialization function $\mathcal{I}$ could speed up convergence, whereby $\mathcal{I}$ can learn and then avoid initializing in uninteresting regions of parameter space. To further test whether different node features could improve the frequency of successful optimization outcomes, we tested several additional node features, sampling the outcome 1000 times each. In total, we tested nine different node feature types: (i) adjacency matrix, (ii) the PCA vectors created from the adjacency matrix, (iii) rate matrix, (iv) the PCA vectors created from the rate matrix, (v) MFPTs (default in other examples), (vi) the PCA vectors created from the MFPT matrix, (vii) eigenvectors of the rate matrix, (viii) trainable embeddings, (ix) a neural network. The feature comparison results (Fig. 4b) show that it is possible to use improved node features, such as the adjacency matrix or a trainable embedding in this case, to achieve a much more successful optimization for the 1D model potential-derived network.
@@
When using the rate matrix as node features in this example, the GNN fails to find the correct clustering even once. This may be due to the difficulty in discriminating nodes at high energies (with low populations) based solely on the rate matrix, as these nodes have similar transitions. The adjacency matrix, however, provides more information on the node location in the global structure of the network, leading to a noticeable advantage. Our GNN implementation allows for simple changes like this to enable defining the node features on a problem-specific basis, therefore giving the method additional viability across a wide range of MCs which vary greatly in structure.
@@
Figure 4: (a) Plot showing the 1D potential used for an example of a linear chain network. Nodes are coloured according to the most optimal 4-state clustering for the system. (b) Violin plots for feature comparisons on the 1D potential network. Each experiment was run 1000 times with each set of features. The corresponding distribution of final $\Delta K/KC$ values is shown, along with a dotted blue line signifying the $\Delta K/KC$ value of the optimal solution for each problem.
@@
We also tested the four GNN architectures on the 1D potential, training 1000 models of each variety using the adjacency matrix as node features. Results are shown in Fig. 5b and described in Table 1. We found that simpler models perform better at finding optimal partitionings over more complex models in this framework. Despite the better average performance the GATv2 and TD models provided, no models but the simplest GNN were able to find the optimal clustering over 1000 runs.
The wide variance of the observed results showcases the poor ability of all models to optimize initial partitionings on this network when started from many regions of parameter space, however, in the smaller models this is mitigated by the smaller solution space explored and training time required. The large magnitude of some solutions is indicative of a partitioning problem where there is only a small region of interest in partitioning space, leading larger models like the GATv2 and TD models to struggle to find this region, possibly due to the more complex and unclear gradients calculated when the node embeddings are more co-dependent and additional resources and required to train the extra parameters. There is a sharp cutoff in the performance of all larger models, with the best value attained being consistently around 806 and the corresponding clustering shown in Fig. 5a. We were unable to pinpoint the exact cause of this lower bound. We also note that in current MSM building strategies only metastable clusters are identified, unless discretized collective variables are introduced. Typically, a k-means clustering is performed that assigns very rarely sampled transition regions to metastable states. Therefore the optimal clustering even from more complex GNNs (Fig. 5a) is able to separate the three main metastable states, which is in line with current practical MSM applications. This problem is particularly challenging for GP objective functions that are not derived from kinetics-based measures, and such measures produce qualitatively incorrect optimal clustering, such as Modularity29. Accordingly, the non-optimal solution of the GNNs represented in Fig. 5a corresponds to a lower DBI value than that of Fig. 4a, which suggests that the DBI measure is also inadequate for this system (Fig. S2). The optimal $\Delta K$ clustering, however, is identical to the solution found by PCCA+ in this example and is consistently found by the PTVC method due to the small size of this network.
@@
Figure 5: (a) The clustering corresponding to the lowest $\Delta K/KC$ values found by all GNN models other than GraphSAGE + LD. (b) The distribution of $\Delta K/KC$ values from training 1000 copies of each neural network architecture on the 1D potential derived network. A logarithmic scale is used to highlight the large differences between the different models. The blue dashed line indicates the value associated with the optimal clustering for this system.
@@
Table 1: Network architecture comparison for the 1D potential derived network. The four network architectures are compared in terms of the mean (standard deviation) and minimum for 1000 experiments, along with the percentage of times the optimal solution was found.
@@
PCCA+ 169.69 ( $\pm$ 0.0) 169.69 N/A
Examples derived from numerical simulations of biomolecular systems often result in noisy derived MCs with dense connections, such as in the pentapeptide case. Here, the optimal PCCA+ clustering does not correspond to the optimal $\Delta K$ solution, with four nodes assigned to different clusters, which are highlighted in Fig. 6a and contribute 7.4% of the overall equilibrium probabilities (Fig. S3). As shown in Fig. 7b and Table 2, again the simplest GraphSAGE and linear decoder method outperforms all other GNN architectures similarly to the 1D example above. The PTVC method does not identify the optimal $\Delta K$ clustering in general when using a random initialization, but has been observed to find the correct solution when given the PCCA+ as an initial clustering. Here, we also compared the optimal solutions from different methods in Fig. 7a by evaluating the DBI score. Interestingly, our optimal KC clustering also corresponds to the lowest DBI score, indicating a better clustering using this measure as well.
As the pentapeptide system is relatively small, significant structural differences cannot be identified within the identified small microstates. Correctly identifying these microstates is an open problem, with significant implications on the dynamics of the MSM 65. We expect more significant differences between states for larger networks that correspond to larger systems, going beyond prototype peptide simulations.